OpenStudy (anonymous):

The perimeter of a rectangle is 72 centimeters. The base is 2 times the height. What are the dimensions of the rectangle

OpenStudy (paki):

the perimeter of rectangle is 2(base + height) let the height be = x and the base be = 3x now perimeter = 2(x + 3x) = 72 x = 8x = 72 x = 72/8 x = 9 height = 9... base = 9 * 3 = 27...

OpenStudy (paki):

sorry... base = 2*9= 18... 3 is by mistaken ...

OpenStudy (texaschic101):

2B + 2H = Perimeter 2B + 2H = 72 B = 2H 2(2H) + 2H = 72 4H + 2H = 72 6H = 72 H = 12 B = 2H B = 2(12) B = 24 check.. 2B + 2H = 72 2(24) + 2(12) = 72 48 + 24 = 72 72 = 72 (correct)